OpenStudy (anonymous):

If : \(\color{blue}{\text{2 + 6 + 10 + 14 + 18 + 22 + 26 + 30 + 34 + 38 = 200}}\), Then, \(\color{red}{\text{Add}}\) any \(\color{cyan}{\text{Five}}\) numbers and get \(\color{green}{100}\)..

OpenStudy (anonymous):

i dont really understand the question:|

OpenStudy (anonymous):

Add any of the 5 numbers from 2, 6, 10, 14, 18, 22, 26, 30, 34, 38 and get the sum as 100.. Like: 38 + 34 + 10 + 18 = 100 (But these are 4 numbers sum only).. Getting @Coolsector

OpenStudy (anonymous):

yes ty

OpenStudy (anonymous):

6+10+22+28+34=100

OpenStudy (anonymous):

oh..... no 28

OpenStudy (anonymous):

Ha ha ha ha... @sauravshakya

OpenStudy (anonymous):

Is there something to do with its pattern.

OpenStudy (anonymous):

Every number is average

OpenStudy (anonymous):

impossible i think

OpenStudy (anonymous):

Just add any 5 numbers, to get result as 100.. I am unable in this...

OpenStudy (anonymous):

@sara12345 , it is impossible only until we don't get the answer for this.. If we get that, then it is possible...

OpenStudy (anonymous):

we can prove its impossible il try

OpenStudy (anonymous):

And I am not typing...

OpenStudy (anonymous):

Have you deleted your comment ?? @hartnn

hartnn (hartnn):

@waterineyes i don't think i will find a single post where you have commented and then NOT written that you were not typing..... yes, i just asked whether we can repeat numbers....

hartnn (hartnn):

getting 90 and 110, but not 100

OpenStudy (anonymous):

We can't repeat numbers @hartnn ..

OpenStudy (anonymous):

Well I can prove it is not possible.

OpenStudy (anonymous):

I can show you that single post if you are looking for the post ..@hartnn

hartnn (hartnn):

no, i am not looking :P haha..

OpenStudy (anonymous):

Oh My God my english...!!!

OpenStudy (anonymous):

Let the numbers five numbers be 2+(n1-1)4 , 2+(n2-1)4 , 2+(n3-1)4 , 2+(n4-1)4 , 2+(n5-1)4 =>4n1-2 , 4n2-2 ,4n3-2 , 4n4-2 , 4n5 -2 Now, For the sum to be 100 4(n1+n2+n3+n4+n5)-10=100 4(n1+n2+n3+n4+n5)=110 Now, 110 is not divisible by 4 Thus, NOT POSSIBLE.

OpenStudy (anonymous):

What! NOBODY got it

hartnn (hartnn):

i got it

OpenStudy (anonymous):

i like ur proof,

OpenStudy (anonymous):

Modified slightly : let the numbers five numbers be 2+4n1 , 2+4n2 , 2+4n3 , 2+4n4 , 2+4n5 Now, For the sum to be 100 4(n1+n2+n3+n4+n5)+10=100 4(n1+n2+n3+n4+n5)=90 Now, 90 is not divisible by 4 Thus, NOT POSSIBLE.

hartnn (hartnn):

see,i told ya, 90 and 110 are possible, not 100...and now they proved it :) nice proof saura :)

OpenStudy (anonymous):

@waterineyes did u get it?

OpenStudy (anonymous):

I am still waiting for someone's reply...

OpenStudy (anonymous):

And yes I got the proof of both of you and I liked it... Nice...

hartnn (hartnn):

i hope he doesn't write , that he is not typing

OpenStudy (anonymous):

that was funny ;p

OpenStudy (anonymous):

I don't think so at this time...

OpenStudy (anonymous):

May be he will come up with 3rd method to prove the above...

OpenStudy (anonymous):

there is only one method so far, saura's !

OpenStudy (anonymous):

And sara's

hartnn (hartnn):

is it only i or everone feels that the names rhyme....sauras and saras....

OpenStudy (anonymous):

phonetically they're diff

hartnn (hartnn):

jhonny! you were writing something ! ? plz feel free to express your opinion, we are waiting....

jhonyy9 (jhonyy9):

2+6+22=30 14+26=40 22+34+14=70 38+10=48 6+14+30=50 38+22=60 38+14+22=74+18+6=98 YES ,THIS IS NOT POSSIBLE FROM 5 NUMBERS SURE

OpenStudy (anonymous):

Okay I agree to all.... Thanks to all for your great time here on this post...

OpenStudy (anonymous):

with 6 its possible

OpenStudy (anonymous):

with 4 it is possible too...

OpenStudy (anonymous):

with 7,8 its not possible

OpenStudy (anonymous):

Yep with six looks possible as 112 0r 88 is divisible by 4

OpenStudy (anonymous):

With 11 it is not possible.. Ha ha ha ha...

OpenStudy (anonymous):

lol... 9 also its not possible

OpenStudy (anonymous):

1 not possible too ...

OpenStudy (anonymous):

thanker water and saura nice exercise :)

OpenStudy (anonymous):

thanks*

OpenStudy (anonymous):

And I am still not typing... Ha ha ha...

OpenStudy (anonymous):

Great thanks to you all... You can help others now if you want...

OpenStudy (anonymous):

0+2+6+10+14+18+22+26+30+34+38=200 so 0+14+22+38+26=100 your question is solve ok

OpenStudy (anonymous):

Yes you can do this but this is bit different from our main question.. Adding 0 will simply solve the problem and there is no doubt in it... I appreciate your thinking @yugraaz ...

OpenStudy (anonymous):

not possible